# Position Paper: Challenges and Opportunities in Topological Deep Learning

### v1

@@ Abstract
Topological deep learning (TDL) is a rapidly evolving field that uses topological features to understand and design deep learning models. This paper posits that TDL may complement graph representation learning and geometric deep learning by incorporating topological concepts, and can thus provide a natural choice for various machine learning settings. To this end, this paper discusses open problems in TDL, ranging from practical benefits to theoretical foundations. For each problem, it outlines potential solutions and future research opportunities. At the same time, this paper serves as an invitation to the scientific community to actively participate in TDL research to unlock the potential of this emerging field.
@@
Figure 1: Topological spaces enrich deep learning methods from a variety of perspectives. (a): A topological space, modeled as a cell complex, enables a flexible molecular representation. Such a representation can enhance the performance of a deep learning model supported on this space (Bodnar et al., 2021a). (b): A topological neural network enables the processing of data via a higher-order message-passing scheme on a topological space, finding a wide range of applications, from computer graphics to drug discovery (Hajij et al., 2023b). (c): Topological spaces allow hierarchical representations of the underlying data that naturally correspond to pooling operations in deep learning (Hajij et al., 2023b). (d): Topological characteristics of the underlying data are crucial when selecting a neural network architecture. Data from a knotted structure in $\mathbb{R}^{3}$, such as the one shown, cannot be embedded in $\mathbb{R}^{2}$ with a single layer multilayer perceptron (MLP) from $\mathbb{R}^{3}$ to $\mathbb{R}^{2}$ (Olah, 2014).
@@ 1 Introduction
Topology is concerned with the study of properties that remain invariant under continuous deformations, and affords a powerful lens through which the global structure of data can be discerned. By characterizing connectivity, continuity, and shape-preserving transformations within datasets, topological tools such as persistent homology (Carlsson, 2009; Edelsbrunner & Harer, 2010) have become powerful methods to capture essential structures and patterns that elude conventional methods. The ability to quantify the topological signatures of data can enhance the robustness of traditional machine learning models, enabling them to discern a greater variety of meaningful structures in diverse and complex datasets. This use of topological ideas, which often runs under the umbrella of topological data analysis (TDA), is by now a well-established field of research. For example, persistent homology has led to victories in the D3R Grand Challenges, a worldwide annual competition series in computer-aided drug design  (Nguyen et al., 2019, 2020).
@@
Olah (2014) contributed one of the first works on TDL, showing that the topology of the underlying space must be considered when choosing a neural network architecture. In the early stages, the term ‘TDL’ was often used to refer to the incorporation of features generated by persistent homology within the input pipeline of a deep neural network (DNN; Cang & Wei, 2017; Hofer et al., 2017). However, this paper uses the term ‘TDL’ to refer to the collection of ideas and methods related to the use of topological concepts in deep learning. Following Hensel et al. (2021), TDL methods can be used in an observational fashion, improving the understanding of existing deep learning models and their topological formulations, or in an interventional fashion, allowing deep learning architectures to treat data supported on higher-order domains, such as cell complexes (Hajij et al., 2020; Bodnar et al., 2022; Giusti et al., 2023), simplicial complexes (Bodnar, 2022; Schaub et al., 2022), sheaves (Hansen & Ghrist, 2019; Bodnar et al., 2021b), combinatorial complexes (Hajij et al., 2023b), and hypergraphs (Feng et al., 2019; Kim et al., 2020; Bai et al., 2021b).
@@
In TDL, multi-way interactions (also known as higher-order relations) between entities constitute useful features capable of embedding topological structure via deep learning algorithms. Higher-order relations capture long-distance or seemingly disparate connections in a system, providing scope for effective or robust message-passing schemes (Hajij et al., 2020; Roddenberry et al., 2021; Bodnar, 2022; Hajij et al., 2023b; Yang & Isufi, 2023). While higher-order relations are at the epicenter of TDL, ongoing research aims to scientifically discern the relative advantages of TDL.
@@
Beyond higher-order relations, the topological view may also capture the regularities inherent to manifolds. Examples include ‘remeshing symmetry’ over manifolds, such as being invariant to different triangulations of a sphere or inducing similar behaviors at different meshing resolutions. These regularities would be perhaps not impossible, but quite challenging, to express using purely combinatorial constructs offered by graph representation learning (GRL), while TDL is a natural domain to think about such effects. Figure 1 shows a range of examples in which topological approaches to deep learning are useful and sometimes crucial.
@@
Position. This paper posits that TDL holds a place in the contemporary landscape of AI research (see figure 1). Higher-order relations and topological invariances are expected to play a crucial role in articulating the position of TDL in relation to overlapping fields, such as GRL or GDL. Finding the conditions under which TDL becomes the go-to machine learning tool for practitioners is one of the main research problems in TDL. This paper builds on the engineering and mathematical knowledge of the community to identify possible research directions to solve this problem.
@@ 2 Examples, Datasets and Benchmarks
This section addresses the need for applications that corroborate the advantages of TDL in certain settings. In addition, it proposes ways to enrich the range of datasets and benchmarks that can facilitate the evaluation of TDL methods.
@@ 2.1 Examples
Compelling applications of TDL are needed in which the topological properties of the data empirically demonstrate a competitive edge.
Several application areas are plausible candidates for TDL to shine, since their underlying domains give rise to topological structures. To encourage the development of TDL examples, the presence or emergence of topological structures in various scientific disciplines is elaborated in appendix B.
@@ 2.2 Datasets
Higher-order data can be collected or synthesized. The TDL application areas of subsection 2.1 are potential candidates to collect higher-order data. Conversely, a graph can be lifted to a higher-order domain (Papillon et al., 2023). Lifting procedures for graphs supply mechanisms for the generation of synthetic topological data. This is why a survey of lifting procedures and associated message-passing schemes would be useful for TDL. Moreover, a generalization of graph-rewiring approaches and analyses to higher-order topological structures would be a fruitful direction.
@@ 3 Software
To the best of the authors’ knowledge, four software packages provide functionality on higher-order structures, namely HyperNetX (Liu et al., 2021), XGI (Landry et al., 2023), DHG (Feng et al., 2019), and TopoX (Hajij et al., 2024). HyperNetX enables computations on hypergraphs, while XGI provides similar functionality on hypegraphs and simplicial complexes. DHG is a deep learning package for graphs and hypergraphs. TopoX is a suite of Python packages designed to compute and learn with topological neural networks. The suite consists of three packages, TopoNetX, TopoEmbedX and TopoModelX. TopoNetX supports computations on graphs and higher-order domains, including colored hypergraphs, simplicial complexes, cell complexes, path complexes and combinatorial complexes. TopoEmbedX provides methods to embed higher-order domains into Euclidean domains. TopoModelX implements the majority of topological neural networks surveyed in Papillon et al. (2023).
@@
However, despite numerous theoretical advances in TDL, practical implementations are scarce due to the limited availability of easy-to-use software packages for deep learning on higher-order structures. TDL is a broad research area, so the contemporary TDL software landscape needs to be enriched with more functionality to cover additional higher-order domains, graph-lifting and rewiring algorithms, TDL models, and learning algorithms, including algorithms for probabilistic inference.
@@
To accelerate the development of TDL software, more human capital and financial investment are required. Since the currently available TDL software packages are open source, there are software engineers and researchers who volunteer to improve these packages. Coding challenges (Papillon et al., 2023) help in this direction. However, more programmers are needed. Academic research grants are one funding avenue for computer scientists to contribute to the development of TDL software. To attract financial resources from investors, it is expected to demonstrate the value of TDL. In other words, solving open problem 1 (compelling examples of TDL) creates conditions to address open problem 4 (resources for the development of TDL software).
@@ 4 Complexity and Scalability
While the development of software for TDL is one problem, another imminent question is whether the benefits of TDL outweigh the resulting costs. To spark some thoughts towards an answer, this section discusses the complexity and scalability implications of TDL.
@@ 4.1 Complexity
A cost-benefit analysis framework for TDL has not been formalized. The time and space complexity of TDL must be factored in systematically to enable informed decision-making when it comes to choosing between TDL, GDL or GRL.
@@
Another approach to dissecting the notion of complexity in TDL is to study the propagation of topological information through neural network layers. For example, Naitzat et al. (2020) have demonstrated that neural networks transform topologically complicated input data into topologically simpler forms as they pass through the layers. Similarly, Wheeler et al. (2021) have studied how the persistence homology of data transforms as the data pass through successive layers of a deep neural network. While Naitzat et al. (2020) and Wheeler et al. (2021) do not focus on input data supported on higher-order structures, such works provide plausible ways of characterizing topological complexity, for instance, via layer-wise transformations of the underlying topology of the data.
@@ 5 Explainability, Generalization and Fairness
This section focuses on the problems of explainability, generalization, and fairness in deep learning. It discusses how TDL can play an important role in solving these problems.
@@ 6 Theoretical Foundations
This section is dedicated to the theoretical foundations of TDL. It focuses on open problems and associated research directions regarding the advantages of TDL, topological representation learning, and transformers in TDL.
@@ 6.1 Advantages of Topological Deep Learning
Why is topology relevant in deep learning? A first step in answering this question is to identify empirically verifiable examples in which TDL outperforms traditional learning methods in some respect (see also open problem 1). Validating such possible empirical findings requires the establishment of theoretical answers. How can theory justify the relevance of topology in deep learning? Can theory prove (or disprove) that higher-order domains provide a richer representation than vector spaces?
When do higher-order relations become useful? Higher-order relations may not benefit every application. That may be the case, especially when higher-order relations are handcrafted, for example, when they are formed by augmenting the input data. In other words, not all higher-order relations are created equal. Is it possible to discover the context in which higher-order relations become useful? Is it possible to pinpoint higher-order relations that lead to improved downstream performance? Once discovered, can these relations be certified as semantically meaningful? Can physical or biological relevance be drawn from higher-order relations with visual tools or semantic metrics? Can higher-order relations help scientists understand phenomena that are currently obscure?
@@
Generative TDL. Generative modeling aims to learn the distribution of training samples by generating new instances. While image generation is now well studied  (Rombach et al., 2022), generating real-world data residing in non-Euclidean or mixed domains presents an unresolved challenge. This has led to research on the generation of 3D point clouds (Zhao et al., 2019; Zeng et al., 2022), meshes (Liu et al., 2023b; Poole et al., 2022), and graphs (Guo & Zhao, 2022; Vignac et al., 2023). Yet, there is no unifying representation that can generate all these 3D representations. This is one of the premises of TDL. In the field of structural biology, the generation of molecules and proteins is currently approached as graph generation (Vignac et al., 2023; Hoogeboom et al., 2022; Ingraham et al., 2019; Zhang et al., 2023) or by generative modeling on 3D rigid bodies (Yim et al., 2023). Generative modeling for graphs generally involves creating new graphs from scratch. However, for many applications where data have inherent geometric structures, relying solely on graphs, as done in Veličković (2023), can potentially omit useful higher-order information. Recent studies have shown that considering higher-order relations yields more accurate representations of both scene graphs and molecular structures (Zhan et al., 2022; Watson et al., 2023). In addition to 2D and 3D computer vision, diffusion models have become the state-of-the-art in graph-based molecular generation (Xu et al., 2022; Vignac et al., 2023; Haefeli et al., 2022; Jing et al., 2023). In particular, Jing et al. (2023) have introduced torsional diffusion, a hybrid method that uses topological principles to diffuse in a toroidal space, effectively narrowing the search space and improving both model performance and inference speed. This approach demonstrates how topology can enhance existing generative methods. Despite these advances, the development of higher-order or topological diffusion and flow-matching models remains an uncharted territory. Data characterized by higher-order structures align naturally with TDL. In some cases, such structures are not inherently present, but can be inferred (Tang et al., 2023a, b), raising questions about the nature and implications of generating higher-order features absent from the original data. The assignment of specific semantic meanings and structures to these generated forms is an ongoing research challenge, paving the way for more interpretable TDL. This area of research, both recent and progressive, signals immense potential for the advancement of generative TDL.

### v3

# Position: Topological Deep Learning is the New Frontier for Relational Learning

@@ Abstract
Topological deep learning (TDL) is a rapidly evolving field that uses topological features to understand and design deep learning models. This paper posits that TDL is the new frontier for relational learning. TDL may complement graph representation learning and geometric deep learning by incorporating topological concepts, and can thus provide a natural choice for various machine learning settings. To this end, this paper discusses open problems in TDL, ranging from practical benefits to theoretical foundations. For each problem, it outlines potential solutions and future research opportunities. At the same time, this paper serves as an invitation to the scientific community to actively participate in TDL research to unlock the potential of this emerging field.
@@
Figure 1: Topological spaces enrich deep learning methods from a variety of perspectives. (a): A topological space, modeled as a cell complex, enables a flexible molecular representation. This representation can improve the performance of a deep learning model supported on this space (Bodnar et al., 2021a). (b): Topological neural networks enable the processing of data, for example via higher-order message-passing schemes on a topological space. These networks find a wide range of applications, from computer graphics to drug discovery (Hajij et al., 2023b). (c): Topological spaces allow hierarchical representations of the underlying data that naturally correspond to pooling operations in deep learning (Hajij et al., 2023b). (d): The topological characteristics of the underlying data are crucial when selecting a neural network architecture. Data from a knotted structure in $\mathbb{R}^{3}$, such as the one shown, cannot be embedded in $\mathbb{R}^{2}$ with a single layer multilayer perceptron (MLP) from $\mathbb{R}^{3}$ to $\mathbb{R}^{2}$ (Olah, 2014).
@@ 1 Introduction
Topology is concerned with the study of properties that remain invariant under continuous deformations, and affords a powerful lens through which the global structure of data can be discerned. By characterizing topological features (including connected components, loops, and voids across multiple scales), topological tools such as persistent homology (Carlsson, 2009; Edelsbrunner & Harer, 2010) have become powerful methods to capture essential structures and patterns that elude conventional methods. The ability to quantify the topological signatures of data can enhance the robustness of traditional machine learning models, enabling them to discern a greater variety of meaningful structures in diverse and complex datasets. This use of topological ideas, which often runs under the umbrella of topological data analysis (TDA), is by now a well-established field of research. For example, persistent homology has led to victories in the D3R Grand Challenges, a worldwide annual competition series in computer-aided drug design  (Nguyen et al., 2019, 2020). Moreover, topological descriptors such as contour trees, merge trees, Reeb graphs, mapper, Morse and Morse-Smale complexes are established tools in scientific data analysis and visualization (Banerjee et al., 2020; Hu et al., 2021b).
@@
Why TDL? To articulate why TDL plays a critical role in the encoding, modeling and analysis of relational data, four practical advantages of TDL are presented. First, the topology of the underlying data space determines the choice of possible neural network architectures. Second, topological domains enable the modeling of data containing multi-way interactions (also known as higher-order relations). Third, TDL captures regularities inherent to manifolds, such as ‘remeshing symmetry’. Fourth, TDL captures topological equivariances in the data. In summary, TDL takes into account topological characteristics that appear in relational data, and therefore is a natural choice for various machine learning problems. Figure 1 shows a range of examples in which topological approaches to deep learning are useful and sometimes crucial.
@@
Choice of neural network architecture. Olah (2014) contributed one of the first works on TDL, showing that the topology of the underlying space must be considered when choosing a neural network architecture. In the early stages, the term ‘TDL’ was often used to refer to the incorporation of features generated by persistent homology within the input pipeline of a deep neural network (DNN; Cang & Wei, 2017; Hofer et al., 2017). However, this paper uses the term ‘TDL’ to refer to the collection of ideas and methods related to the use of topological concepts in deep learning. Following Hensel et al. (2021), TDL methods can be used in an observational fashion, improving the understanding of existing deep learning models and their topological formulations, or in an interventional fashion, allowing deep learning architectures to treat data supported on higher-order domains, such as cell complexes (Hajij et al., 2020; Bodnar et al., 2022; Giusti et al., 2023), simplicial complexes (Bodnar, 2022; Schaub et al., 2022; Mitchell et al., 2024), sheaves (Hansen & Ghrist, 2019; Bodnar et al., 2021b), combinatorial complexes (Hajij et al., 2023b), and hypergraphs (Feng et al., 2019; Kim et al., 2020; Bai et al., 2021b; Behrouz et al., 2023).
@@
Multi-way interactions. In TDL, multi-way interactions between entities constitute useful features capable of embedding topological structure via deep learning algorithms. Higher-order relations capture long-distance or seemingly disparate connections in a system, providing scope for effective or robust message-passing schemes (Hajij et al., 2020; Roddenberry et al., 2021; Bodnar, 2022; Hajij et al., 2023b; Yang & Isufi, 2023). While graph-based equivariant neural networks have been developed to incorporate certain multi-way interactions (Batatia et al., 2022; Musaelian et al., 2023), many types of multi-way interactions can arise in practice. TDL operates on various topological domains, such as simplicial and cellular complexes, and therefore provides a framework to model a multitude of possible types of multi-way interactions appearing in relational data.
@@
Regularities inherent to manifolds. Beyond higher-order relations, the topological view may also capture the regularities inherent to manifolds. Examples include ‘remeshing symmetry’ over manifolds, such as being invariant to different triangulations of a sphere or inducing similar behaviors at different meshing resolutions. These regularities would be perhaps not impossible, but quite challenging, to express using purely combinatorial constructs offered by graph representation learning (GRL), while TDL is a natural domain to think about such effects.
Topological equivariances. TDL is a natural approach to capture ‘topological equivariances’. For example, if a classification algorithm is meant to identify different knots, then it is useful to understand the stabilizer group involving isotopies of the complement. In general, GNNs and GDL are based on ‘standard groups’. For instance, GNNs adopt the permutation group, and applications of GDL in molecular modeling use subgroups of the Euclidean group, such as the special Euclidean group $\mbox{SE}(n)$ or the special orthogonal group $\mbox{SO}(n)$. TDL incoporates more complex homeomorphism groups that act on a space, while trying to preserve some embedding information. This embedding information can include, for example, the arrangement of critical points in Morse-Smale complexes (Catanzaro et al., 2020) or the nesting of circles which leads to the concept of ‘tree of shapes’ in image processing (Caselles & Monasse, 2009). In particular, image segmentation provides an example of a topological structure that needs to be preserved under homeomorphisms of the domain.
Position. This position paper argues that TDL is the new frontier for relational learning. Relational data constitute a main modality of data emerging from natural and artificial systems (Veličković, 2023). In such systems, sets of of objects are interconnected via binary or higher-order relations, and relational data encode features of these interconnections. TDL provides a framework for learning from binary or higher-order relational data.
While TDL is pivotal in relational learning, it is more generally a valuable framework with many functions in the contemporary AI landscape. TDL informs the choice of neural network architecture subject to the topology of the underlying data space, captures regularities inherent to manifolds, and captures topological equivariances.
@@
Topological concepts are expected to play a crucial role in articulating the position of TDL in relation to overlapping fields, such as GRL or GDL. Finding the conditions under which TDL becomes the go-to machine learning tool for practitioners is one of the main research problems in TDL. This paper builds on the engineering and mathematical knowledge of the community to identify possible research directions to solve this problem.
@@ 2 Examples, Datasets and Benchmarks
This section highlights compelling applications and success stories of TDL in machine learning challenges to corroborate the advantages of TDL in various settings. In addition, it proposes ways to enrich the range of datasets and benchmarks that can facilitate the evaluation of TDL methods.
@@ 2.1 Examples
Compelling applications of TDL have been developed, in which the topological properties of the data empirically demonstrate a competitive edge. However, a broader adaptation of TDL, which can revolutionize relational learning, has not yet taken place in real-world applications.
Several application areas are plausible candidates for TDL to shine, since their underlying domains give rise to topological structures. To encourage the development and adaptation of TDL in practice, the presence or emergence of TDL applications in numerous scientific disciplines is elaborated further in appendix B.
@@ 2.2 Datasets
Higher-order data can be collected or synthesized. The TDL application areas of subsection 2.1 are potential candidates to collect higher-order data. Conversely, a graph can be lifted to a higher-order domain (Papillon et al., 2023a). Lifting procedures for graphs supply mechanisms for the generation of synthetic topological data. This is why a survey of lifting procedures and associated message-passing schemes would be useful for TDL. Moreover, a generalization of graph-rewiring approaches and analyses to higher-order topological structures would be a fruitful direction.
@@ 3 Software
To the best of the authors’ knowledge, four software packages provide functionality on higher-order structures, namely HyperNetX (Liu et al., 2021), XGI (Landry et al., 2023), DHG (Feng et al., 2019), and TopoX (Hajij et al., 2024). HyperNetX enables computations on hypergraphs, while XGI provides similar functionality on hypegraphs and simplicial complexes. DHG is a deep learning package for graphs and hypergraphs. TopoX is a suite of Python packages designed to compute and learn with topological neural networks. The suite consists of three packages, TopoNetX, TopoEmbedX and TopoModelX. TopoNetX supports computations on graphs and higher-order domains, including colored hypergraphs, simplicial complexes, cell complexes, path complexes and combinatorial complexes. TopoEmbedX provides methods to embed higher-order domains into Euclidean domains. TopoModelX implements the majority of topological neural networks surveyed in Papillon et al. (2023a).
@@
However, despite numerous theoretical advances in TDL, practical implementations are scarce due to the limited availability of easy-to-use software packages for deep learning on higher-order structures. TDL is a broad research area, so the contemporary TDL software landscape needs to be enriched with more functionality to cover additional higher-order domains (including domains related to higher-order multiplex networks, and dynamic versions of existing higher-order domains), graph-lifting and rewiring algorithms, TDL models, and learning algorithms (Alain et al., 2023).
@@
To accelerate the development of TDL software, more human capital and financial investment are required. Since the currently available TDL software packages are open source, there are software engineers and researchers who volunteer to improve these packages. Coding challenges (Papillon et al., 2023a) help in this direction. However, more programmers are needed. Academic research grants are one funding avenue for computer scientists to contribute to the development of TDL software. To attract financial resources from investors, it is expected to demonstrate the value of TDL. In other words, solving open problem 1 (compelling examples of TDL) creates conditions to address open problem 4 (resources for the development of TDL software).
@@ 4 Complexity and Scalability
While the development of software for TDL is one problem, another imminent question is whether the benefits of TDL outweigh the resulting costs. The innovative design of TDL architectures can make a difference in practical applications. In the context of such applications, computational complexity and scalability are major challenges that become particularly prominent in TDL due to the inclusion of higher-order data. This section discusses the complexity and scalability implications of TDL.
@@ 4.1 Complexity
A cost-benefit analysis framework for TDL has not been formalized. The increased time and space complexity of TDL relatively to GDL and GRL (due to the inclusion of higher-order data) must be systematically factored in to enable informed decision-making when it comes to choosing between TDL, GDL or GRL.
@@
Another approach to dissecting the notion of complexity in TDL is to study the propagation of topological information through neural network layers. For example, Naitzat et al. (2020) have demonstrated that neural networks transform topologically complicated input data into topologically simpler forms as they pass through the layers. Petri & Leitão (2020) have expanded the work of Naitzat et al. (2020) by showing that it is not necessarily the topology of the data that is simplified, but the topology of the decision boundary. Wheeler et al. (2021) have studied how the persistence homology of data transforms as the data pass through successive layers of a deep neural network. While Naitzat et al. (2020), Petri & Leitão (2020) and Wheeler et al. (2021) do not focus on input data supported on higher-order structures, such works provide plausible ways of characterizing topological complexity, for instance, via layer-wise transformations of the underlying topology of the data.
@@ 5 Explainability, Generalization and Fairness
This section focuses on the problems of explainability, generalization, and fairness. Such problems are common in machine learning and have been explored in other contexts, but how these can get translated in the context of TDL is largely unexplored. This section discusses how TDL can play an important role in solving these problems.
@@ 6 Theoretical Foundations
This section is dedicated to the theoretical foundations of TDL. It focuses on open problems and associated research directions regarding the advantages of TDL, topological representation learning (TRL), and transformers in TDL.
@@ 6.1 Advantages of Topological Deep Learning
Why is topology relevant in deep learning? A first step in answering this question has been taken by developing empirically verifiable examples in which TDL outperforms traditional learning methods in some respect (see also open problem 1). Validating such possible empirical findings requires the establishment of theoretical answers. How can theory justify the relevance of topology in deep learning? Can theory prove (or disprove) that higher-order domains provide a richer representation than vector spaces?
When do higher-order relations become useful? Higher-order relations may not benefit every application. Is it possible to discover the context in which higher-order relations become useful (Battiston et al., 2020; Benson et al., 2021; Bick et al., 2023)? Is it possible to pinpoint higher-order relations that lead to improved downstream performance? Once discovered, can these relations be certified as semantically meaningful? Can physical or biological relevance be drawn from higher-order relations with visual tools or semantic metrics? Can higher-order relations help scientists understand phenomena that are currently obscure?
What are the advantages of TDL over GRL? One question is whether in certain cases higher-order relations are not required and the same performance might be achieved by stacking more layers in regular GNNs or by using graph transformer networks (Yun et al., 2019). This can potentially happen, but it is known that neural architecture design matters (especially in the low-data regime) with respect to capturing higher-order relations. For example, the work of Sanford et al. (2023) implies that existing transformers are efficient in modeling pairwise relations, but not three-way relations. Furthermore, TDL may help avoid some of the common pathologies that arise with deeper GNNs, such as vanishing gradients and degradation.
@@
Generative TDL. Generative modeling aims to learn the distribution of training samples by generating new instances. While image generation is now well-studied  (Rombach et al., 2022), generating real-world data supported on non-Euclidean or mixed domains presents an unresolved challenge. This has led to research on the generation of 3D point clouds (Zhao et al., 2019; Zeng et al., 2022), meshes (Liu et al., 2023b; Poole et al., 2022), and graphs (Guo & Zhao, 2022; Vignac et al., 2023). Yet, there is no unifying representation that can generate all these 3D representations. This is one of the premises of TDL. In the field of structural biology, the generation of molecules and proteins is currently approached as graph generation (Vignac et al., 2023; Hoogeboom et al., 2022; Ingraham et al., 2019; Zhang et al., 2023) or by generative modeling on 3D rigid bodies (Yim et al., 2023). Generative modeling for graphs generally involves creating new graphs from scratch. However, for many applications where data have inherent geometric structures, relying solely on graphs, as done in Veličković (2023), can potentially omit useful higher-order information. Recent studies have shown that considering higher-order relations yields more accurate representations of both scene graphs and molecular structures (Zhan et al., 2022; Watson et al., 2023). In addition to 2D and 3D computer vision, diffusion models have become the state-of-the-art in graph-based molecular generation (Xu et al., 2022; Vignac et al., 2023; Haefeli et al., 2022; Jing et al., 2023). In particular, Jing et al. (2023) have introduced torsional diffusion, a hybrid method that uses topological principles to diffuse in a toroidal space, narrowing the search space and improving both model performance and inference speed. This approach demonstrates how topology can enhance existing generative methods. Despite these advances, the development of topological diffusion and flow-matching models remains an uncharted territory. Data characterized by higher-order structures align naturally with TDL. In some cases, such structures are not inherently present, but can be inferred (Tang et al., 2023a, b), raising questions about the nature and implications of generating higher-order features absent from the original data. The assignment of specific semantic meanings and structures to these generated forms is an ongoing research challenge, paving the way for more interpretable TDL. This area of research, both recent and progressive, holds the potential to advance generative TDL.